MIDC News
Year of 2023
September 14, 2023
- 15:16 MST: All SRRL BMS Diffuse radiometers (CM22 #1, CM22 #2, SR25, 8-48, SB 8-48) were swapped out between approx. 10:07 and 10:39 MST today with instruments calibrated in BORCAL 2023-02. With the exception of the SB 8-48, no data loss occurred because automatic substitution was enabled with radiometers that were not being swapped at that particular moment. Logger was updated with new factors at 10:20 MST (CM22 #2, 8-48) and 10:38 MST (CM22 #1, SR25, SB 8-48). -Afshin Andreas
- 15:40 MST: The SRRL BMS Upwelling pyranometers (CMP11, CM3, LI200) were swapped out today with radiometers calibrated in BORCAL 2023-02, factors updated in logger @ 12:07 MST today (CM3, LI200) and 12:13 MST today (CMP11). -Afshin Andreas
- 15:38 MST: The SRRL BMS 1-Axis pyranometers (CMP22, SPLite2, LI200R) were swapped out today with radiometers calibrated in BORCAL 2023-04, factors updated in logger @ 6:44 MST today. -Afshin Andreas
- 15:03 MST: All SRRL BMS Direct Normal Shortwave Radiometers were calibrated in BORCAL SW 2023-04 (via signal patching), this includes: CHP1 #1, CHP1 #2, sNIP, NIP, DR02, MS-57, LI-201. Logger was updated with new calibration factors at 14:50 MST today. -Afshin Andreas
- 15:01 MST: All SRRL BMS Global Horizontal Shortwave Radiometers were calibrated in BORCAL SW 2023-02 (via signal patching), this includes: CMP22 (vent), CMP22 #1, CMP22 #2, CMP11, CM6b, CM3, SR25, MS-80, MS-802, MS-410, MS-602, SPP, PSP, GPP, SP-510, TSP-700, SP-Lite2, ML-01, LI-200R, LI-200, SP-110. Logger was updated with new calibration factors (Z@45) at 08:24 MST on July 3 (16:14 MST on July 3 for CFvsZ for PSP). -Afshin Andreas
- 10:03 MST: There was a SRRL power outage on August 6 from 7:12 to 12:06 MST due to an electrical systems failure at NREL. Various BMS instrumentation that do not have backup power will be down, Data Quality Statements (DQS) will be created later this week for affected instrumentation. -Afshin Andreas
- 10:00 MST: The public MIDC web server was migrated to a new server on June 21, 2023. The new server is much faster. -Afshin Andreas
- 11:27 MST: The SRRL BMS 40-South pyranometers (CMP22, SPLite2, LI200R) were swapped out on 6-28-2023 with radiometers calibrated in BORCAL 2023-03, factors updated in logger @ 14:38 MST on the same day. -Afshin Andreas
- 12:49 MST: The SRRL BMS 2-Axis (global normal) pyranometers (CMP22, SPLite2, LI200R) were swapped out on 6-21-2023 with radiometers calibrated in BORCAL 2023-02, factors updated in logger @ 14:43 MST on the same day. -Afshin Andreas
- 09:26 MST: The SRRL BMS Temp/RH HMP-60 Probe (Data Lab) was swapped out with a calibrated sensor at 09:17 MST today. -Afshin Andreas
- 14:42 MST: The SRRL BMS EKO MS-711 horizontal and the MS-700 upwelling spectroradiometers were removed from service today for calibration against NIST FEL F724. They will report missing value during these times:
- Horizontal MS-711: 8:41-12:58 MST
- Upwelling MS-710: 10:34-13:43 MST
- 17:55 MST: FlatIrons M2 EFM is operational as of 17:39 MST today. -Afshin Andreas
- 10:05 MST: On April 6, an EKO MS-700 spectroradiometer (same instrument that was removed from OTF building) was mounted on the SRRL Radiometer Tower measuring upwelling spectral irradiance. -Afshin Andreas
- 14:26 MST: January through February 2023 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 14:50 MST: The SRRL BMS Temp/RH HMP-60 Probe (SE) was swapped out with a calibrated sensor at 09:12 MST today. -Afshin Andreas
- 17:42 MST: The Flatirons M2 tower was dropped yesterday from 07:36 to 11:24 MST, so that the temperature probes, wind speed and direction sensors could replaced. Calibration constants for new temperature and wind speed sensors were updated in logger program at 09:23 MST yesterday. -Afshin Andreas
- 22:12 MST: The SRRL EKO MS-711, which is part of 2-axis WISER spectroradiometer was replaced with an EKO loaner instrument at 13:39 MST so that NREL's instrument can be shipped for repair. -Afshin Andreas
- 13:44 MST: he SRRL BMS Temp/RH HMP-60 Probe (Deck) was swapped out with a calibrated sensor at 13:08 MST today. -Afshin Andreas
- 11:17 MST: The SRRL BMS Temp/RH HMP-60 Probe (Tower) was swapped out with a calibrated sensor at 9:17 MST today. -Afshin Andreas
- 11:28 MST: Flatirons M2 radiometers were swapped out with instruments calibrated in BORCAL 2022-02 between 10:36 and 11:11 MST today. The Direct signal was also repaired (cable was broke). Logger updated with new factors at 10:54 MST. -Afshin Andreas
- 08:54 MST: The SRRL EKO MS-711 (part of 2-axis WISER Spectroradiometer) has been intermittently failing (reading near zero). Problem started near 15-Oct-2022 and has been progressively getting worse. -Afshin Andreas
- 13:18 MST: SolarTAC PWV data is still unavailable until external agency data processor updates IP address for new 4G cell modem. -Afshin Andreas
- 13:16 MST: Cell modem at SolarTAC has been replaced and data collection has resumed. In addition, Global, Direct, and Diffuse radiometers were swapped out between 11:18 and 12:06 MST today with radiometers calibrated in BORCAL 2022-02. -Afshin Andreas
- 14:17 MST: The SRRL BMS Global 90 PSP and LI-200 Sensors were replaced with sensors calibrated in BORCAL 2022-03&04 between 13:54 and 14:07 MST today. Logger was updated with new factors at 14:05 MST today. -Afshin Andreas
- 14:06 MST: October through December 2022 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 10:36 MST: SRRL AOD ESR-L2 and MRI-L2 data sets are not available since the end of October 2022, because SkyNET may make a revision to the data set. However, the ESR-L2A, MRI-L2A, and SR-CEReS are still being generated. -Afshin Andreas
- 08:42 MST: SRRL BMS pressure sensor is operational. No problem was found (but sensor was power cycled during troubleshooting). -Afshin Andreas
- 14:28 MST: The SRRL BMS pressure sensor failed at 13:15 MST today. It could be due to a clogged pressure port. Until the sensor can be investigated tomorrow, BMS is reporting a fixed value of 813.6 mBar (present value as reported by the SRRL BORCAL reference pressure). -Afshin Andreas
- 11:24 MST: Due to SRRL staff shortage, BMS Maintenace (cleaning) was not performed between December 22 through January 2 (except December 27), data during that time should be used with caution. -Afshin Andreas
- 11:22 MST: Data Collection from SolarTAC has been down since December 23 at 20:11 MST due to the Verizon 3G service shutdown. Prior to the holidays we were unable to obtain a working 4G cell modem, data will be unavailable until a 4G cell modem is obtained. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]